Paul Coddington

Northeast Parallel Architectures Center 111 College Place, 3-228 CST Syracuse University Syracuse, NY 13244

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

November 29, 1993

James Mitchie Power Technologies Inc. 1482 Erie Boulevard Schenectady, NY 12301

Dear Mr. Mitchie,

Here is a copy of the proposal for our research project with Niagara Mohawk Power Corporation. It would be extremely valuable for our work if you could provide us with the following information:

1. Some documentation on the algorithms used in the PTI transient stability program.

2. A copy of any reports on previous studies of using parallel processing for this program.

3. Some idea of the relative times required for solving the generator and network equations.

We would also greatly appreciate access to the actual code at some point, if this could be arranged.

A major part of our project is concerned with investigating the possibility of developing an on-line stability analysis program. For this to be feasible, the analysis of the output of the program would have to be done on-line by an expert system, rather than off-line by an engineer. Investigating the feasibility of developing such an expert system is another part of our project. GE has a similar project, utilizing their ENGINEOUS expert systems program and the PTI stability program. We will be meeting with representatives of the GE project soon to discuss a possible collaboration, on both the expert systems and parallel processing aspects of the problem.

I would like to talk to you about this aspect of our work after our meeting with GE. In order to implement a prototype expert system, access to the PTI program (which we currently do not have) would be invaluable.

Yours sincerely,